Solve the equation for x in the interval 0≤x ≤2 π:
2sin²x-sinx-1=0

Answer by stanbon(48558)

Solve the equation for x in the interval 0≤x ≤2 π:
2sin²x-sinx-1=0
Factor to get:
(2sinx+1)(sinx-1) = 0
sinx = -1/2 or sinx = 1
x = (11/6)pi or (7/6)pi or pi/4 or (3/4)pi
==============================================
Cheers,
Stan H.